# Diffusion-Based Image-to-Brain Signal Generation with Cross-Attention Mechanisms for Visual Prostheses

### v1

@@ Abstract
Visual prostheses have shown great potential in restoring vision for blind individuals. On the one hand, researchers have been continuously improving the brain decoding framework of visual prostheses by leveraging the powerful image generation capabilities of diffusion models. On the other hand, the brain encoding stage of visual prostheses struggles to generate brain signals with sufficient biological similarity. Although existing works have recognized this problem, the quality of predicted stimuli still remains a critical issue, as existing approaches typically lack supervised signals from real brain responses to validate the biological plausibility of predicted stimuli. To address this issue, we propose a novel image-to-brain framework based on denoising diffusion probabilistic models (DDPMs) enhanced with cross-attention mechanisms. Our framework consists of two key architectural components: a pre-trained CLIP visual encoder that extracts rich semantic representations from input images, and a cross-attention enhanced U-Net diffusion model that learns to reconstruct biologically plausible brain signals through iterative denoising. Unlike conventional generative models that rely on simple concatenation for conditioning, our cross-attention modules enable dynamic interaction between visual features and brain signal representations, facilitating fine-grained alignment during the generation process. We evaluate our framework on two multimodal datasets (THINGS-EEG2 and THINGS-MEG) to demonstrate its effectiveness in generating biologically plausible brain signals. Moreover, we visualize the training and test M/EEG topographies for all subjects on both datasets to intuitively demonstrate the intra-subject variations and inter-subject variations in M/EEG signals.
@@ 1 Introduction
Visual prostheses are advanced medical devices designed to restore partial vision for individuals with severe visual impairments or blindness, often caused by conditions such as retinitis pigmentosa (RP) and age-related macular degeneration (AMD) (Zrenner, 2013; Busskamp & Roska, 2011). These devices use an image sensor to capture external visual scenes and a processing framework to predict stimuli for an implanted electrode array (Humayun et al., 2012; Goetz & Palanker, 2016; Soltan et al., 2018) (we call this process brain encoding). The electrode array stimulates ganglion cells with the predicted stimuli, evoking visual perception (a pattern of localized light flashes, visual percept, or ’phosphene’) in the retina (van der Grinten et al., 2024; Blom et al., 2010; Berry et al., 2017; Sahel et al., 2021; Granley et al., 2023) (this process is also referred to as brain decoding (Benchetrit et al., 2023)). The framework of visual prostheses is illustrated in Figure 1.
In the past few years, brain decoding has made significant progress (Lin et al., 2022; Scotti et al., 2023; Wang et al., 2024b; Li et al., 2024). Specifically, Mind-Reader (Lin et al., 2022), MindEye (Scotti et al., 2023), and MindBridge (Wang et al., 2024b) utilize the high spatial resolution of functional magnetic resonance imaging (fMRI) to generate phosphenes. However, due to the high cost and low temporal resolution of fMRI limiting their applications in brain-computer interfaces (BCIs), Li et al. (Li et al., 2024) not only leverage the high temporal resolution of electroencephalography (EEG) signals to evoke visual percepts, but also demonstrate the versatility of their work on magnetoencephalography (MEG) signals. More importantly, these studies (Lin et al., 2022; Scotti et al., 2023; Wang et al., 2024b; Li et al., 2024) utilize multimodal datasets (Allen et al., 2022; Gifford et al., 2022) that include not only brain signals but also image data. Therefore, when training models, whether brain signals or image data are required, corresponding supervised signals can be provided to validate the model’s output.
@@
Figure 1: The framework of the visual prostheses. Visual prostheses utilize an image sensor to capture natural scenes. A processing framework takes the recorded signals as input and predicts the stimuli for the retinal prosthesis. A phosphene model receives stimulation from the implanted prosthesis and evokes visual perception (phosphenes). The performance of the framework is evaluated by comparing the similarity between the original image and the visual perception.
To address the aforementioned issues, we propose an innovative image-to-brain framework. We employ a denoising diffusion probabilistic model (DDPM) (Ho et al., 2020) enhanced with cross-attention mechanisms. This framework consists of two core components: a CLIP visual encoder and a cross-attention enhanced U-Net diffusion model. The CLIP visual encoder extracts rich semantic representations from input images using pre-trained Vision Transformer (ViT-L/14) (Radford et al., 2021). The U-Net diffusion model reconstructs brain signals through iterative denoising, while cross-attention mechanisms enable fine-grained alignment between visual features and brain signal representations during the generation process. To validate our method’s effectiveness, we conduct experiments on multimodal datasets containing both brain signals and image data. With these datasets, we can directly learn the mapping from images to brain signals using the ground truth brain responses as supervision signals.
@@
  • We propose a novel image-to-brain framework based on diffusion models that achieves high-quality conversion from images to brain signals, advancing the technical foundation for visual prostheses.
@@
  • We validate our method through comprehensive experiments and plot the training and test M/EEG topographies for each subject on both datasets to intuitively demonstrate the intra-subject variations and inter-subject variations of M/EEG signals.
@@ 2 Related Works
In recent years, brain decoding has made significant progress by leveraging the powerful generative capabilities of diffusion models (Lin et al., 2022; Scotti et al., 2023; Wang et al., 2024b; Li et al., 2024; Xu et al., 2023). In contrast, the development of brain encoding has progressed relatively slowly. Despite ongoing research efforts to improve the quality of predicted stimuli (Granley et al., 2022; 2023; Wang et al., 2024a), these studies fail to utilize real stimuli as supervised signals for evaluating the biological similarity of predicted stimuli, thereby limiting the vision restoration efficacy of visual prostheses to a low level (Montazeri et al., 2019).
To address the issues mentioned above, we use brain signals from multimodal datasets (THINGS-EEG2, THINGS-MEG) as supervised signals to improve the biological similarity of predicted stimuli, thereby refining the image-to-brain framework.
@@
To understand the complex interplay between visual stimuli and human brain activity, Lu et al. (Lu & Golomb, 2024) propose the Img2EEG method, which inputs images into deep learning modules to generate EEG signals. This work highlights the importance of establishing reliable mappings between visual inputs and corresponding brain responses, which is crucial for advancing visual prostheses applications.
@@ 3.3.2 Cross-Attention Mechanism
The key innovation of our approach lies in the integration of cross-attention mechanisms within the U-Net architecture. Unlike conventional generative models that use simple concatenation or addition for conditioning, our cross-attention modules (Vaswani et al., 2017) enable fine-grained alignment between visual features and brain signals. We modify the standard U-Net by incorporating cross-attention blocks in both the encoder and decoder paths. These blocks enable dynamic interaction between the brain signal representations and visual features during the denoising process. For each cross-attention layer, given the intermediate brain signal representation $\mathbf{H}_{\text{brain}}$ and visual embedding $\mathbf{h}_{\text{img}}$, the cross-attention is computed as:
@@ 3.5 Test Stage
We start with pure Gaussian noise $\mathbf{y}_{T}\sim\mathcal{N}(0,\mathbf{I})$ and iteratively denoise it using the learned reverse process. Given an input image, we first extract visual features:
@@
Then, for each time step $t=T,T-1,\ldots,1$, we perform the denoising step:
$\mathbf{y}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{y}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\bm{\epsilon}_{\theta}(\mathbf{y}_{t},t,\mathbf{h}_{\text{img}})\right)+\sigma_{t}\mathbf{z},$ (14)
where $\mathbf{z}\sim\mathcal{N}(0,\mathbf{I})$ and $\sigma_{t}$ is the noise variance at step $t$. The final output $\mathbf{y}_{0}$ represents the generated brain signal corresponding to the input image.
@@ 4.2 Experiment Details
During training, we use the complete framework shown in Figure 2 with paired image-brain signal data. For testing, we perform image-to-brain signal conversion by first extracting image embeddings via the CLIP visual encoder, then feeding these embeddings to the trained U-Net decoder to generate brain signals through iterative denoising.
Model Architecture: Our diffusion model employs a UNet2DConditionModel from the Hugging Face diffusers library (von Platen et al., 2022) with dataset-specific configurations. For EEG signals, we use a sample size of $(63,250)$ with 63 channels and 250 temporal sampling points. For MEG signals, we configure the model with a sample size of $(271,200)$ with 271 channels and 200 temporal sampling points. Both models use 1 input and output channel, 4 downsampling and upsampling blocks with channel dimensions of $(128,256,512,512)$. The downsampling path consists of: two DownBlock2D layers followed by two CrossAttnDownBlock2D layers, while the upsampling path includes: two CrossAttnUpBlock2D layers followed by two UpBlock2D layers. Cross-attention mechanisms are specifically integrated in the deeper layers to enable fine-grained alignment between visual features and brain signal representations (Ablation study on cross-modal learning can be seen in Appendix A.1). The cross-attention dimension is set to 768 to match the CLIP embedding dimension. We use the ViT-L/14 variant of CLIP as our visual encoder to extract rich semantic representations from input images.
@@ 4.3 Performance Evaluation
Within-subject Performance: Tables 1 and 3 present the within-subject results, where models are trained and tested on data from the same subject. For the THINGS-EEG2 dataset, our method achieves an average MSE of 0.217 and PCC of 0.188 across 10 subjects. The performance varies across subjects, with Subject 1 achieving the best MSE (0.178) and Subject 8 showing the highest PCC (0.231). For the THINGS-MEG dataset with 4 subjects, we obtain an average MSE of 0.763 and PCC of 0.122. Among all subjects, Subject 1 achieves the lowest MSE of 0.607, while Subject 2 obtains the highest PCC of 0.198.
Cross-subject Generalization: The cross-subject generalization results provide insights into the transferability of our framework across different subjects. From Table 2, we observe that each trained source model demonstrates consistent performance across other subjects, with very small standard deviations (ranging from 0.003 to 0.007), indicating good cross-subject generalization capability of our method. In contrast, Table 4 shows that the source mean and source standard deviation reveal larger performance fluctuations when source models are tested on different target subjects. However, examining the target mean and target standard deviation, we find that different source models achieve similar performance on the same target subject, suggesting that each subject in the THINGS-MEG dataset exhibits inherent differences. The detailed cross-subject PCC results are presented in Table 9 and Table 10 (See Appendix A.2).
@@ 4.4 Visualization Analysis
I have plotted the topography for all subjects in both the THINGS-EEG2 and THINGS-MEG datasets, as detailed in Figures 3, 4, 5, 6, 7, and 8 (See Appendix A for Figures 5, 6, 7, and 8).
@@
5 Discussion and Conclusion
The work most similar to ours is (Lu & Golomb, 2024). However, while Lu et al. also use the THINGS-EEG2 dataset in their paper, their framework outputs EEG signals with dimensions of [17 channels, 50 time points], whereas the original EEG signals in THINGS-EEG2 have dimensions of [63 channels, 250 time points]. This means that Lu et al.’s work only establishes correspondence between images and a subset of the corresponding EEG signals, failing to capture the complete brain dynamics during object recognition. In contrast, our method maintains the same output dimensions as the original THINGS-EEG2 dataset. Furthermore, we also conduct experiments on MEG signals, which have higher spatial resolution than EEG signals, to demonstrate the versatility of our approach.
Of course, our work also has limitations. Due to the substantial computational overhead introduced by diffusion models and cross-attention mechanisms, combined with constraints from our hardware equipment, we are unable to employ deeper and more complex neural network architectures to achieve better model performance.
In conclusion, we propose an image-to-brain framework (we call image-to-brain process as brain encoding). This framework uses diffusion models to complete the brain signal reconstruction task and uses cross-attention to achieve alignment between the two modalities of images and brain signals. We conduct experiments on both THINGS-EEG2 and THINGS-MEG datasets, demonstrating the compatibility of our framework with both EEG and MEG signals. Meanwhile, we also plot topographies of EEG and MEG signals, allowing us to more intuitively observe the conditions of these two datasets.
@@ A.1 Ablation Study on Cross-Modal Learning
Tables 5 and 7 present the MSE results for the THINGS-EEG2 and THINGS-MEG datasets, respectively. Our cross-attention approach achieves the lowest average MSE on both datasets (0.217 for EEG and 0.763 for MEG), demonstrating superior performance compared to concatenation (0.228 for EEG and 0.808 for MEG) and addition (0.227 for EEG and 0.811 for MEG) methods.
Tables 6 and 8 present the results on the THINGS-EEG2 and THINGS-MEG datasets, respectively, with all values reported as Pearson Correlation Coefficient (PCC). The cross-attention mechanism achieves the highest average PCC on both datasets (0.188 for EEG and 0.122 for MEG), outperforming both concatenation (0.180 for EEG and 0.109 for MEG) and addition (0.172 for EEG and 0.118 for MEG) methods. The consistent advantage of cross-attention across both datasets and across different evaluation metrics indicates that explicitly modeling the dynamic interactions between brain signals and visual features leads to better performance in cross-modal learning tasks.
A.2 Cross-Subject PCC Results
Table 9: Cross-subject PCC results on THINGS-EEG2 dataset
Train Subject Test Subject Source Stats
1 2 3 4 5 6 7 8 9 10 Mean Std
1 0.145 0.132 0.152 0.128 0.106 0.161 0.151 0.084 0.163 0.136 0.025
2 0.107 0.078 0.141 0.088 0.096 0.099 0.125 0.072 0.122 0.103 0.024
3 0.115 0.096 0.120 0.125 0.086 0.070 0.096 0.053 0.134 0.099 0.027
4 0.112 0.121 0.095 0.110 0.087 0.113 0.125 0.094 0.108 0.107 0.012
5 0.079 0.087 0.082 0.105 0.080 0.071 0.111 0.058 0.090 0.084 0.017
6 0.066 0.077 0.059 0.098 0.091 0.078 0.106 0.078 0.098 0.083 0.016
7 0.123 0.099 0.051 0.134 0.091 0.086 0.106 0.108 0.094 0.099 0.024
8 0.110 0.129 0.089 0.142 0.137 0.131 0.119 0.076 0.125 0.117 0.023
9 0.067 0.081 0.038 0.107 0.071 0.074 0.101 0.069 0.070 0.075 0.020
10 0.140 0.146 0.140 0.132 0.122 0.119 0.112 0.132 0.077 0.124 0.021
Table 10: Cross-subject PCC results on THINGS-MEG dataset
Train Subject Test Subject Source Stats
1 2 3 4 Mean Std
1 0.080 0.038 0.069 0.062 0.021
2 0.088 0.075 0.051 0.071 0.019
3 0.055 0.071 0.045 0.057 0.013
4 0.027 0.033 0.020 0.027 0.007
Target Mean 0.057 0.061 0.044 0.055 0.054
Target Std 0.031 0.024 0.028 0.012 0.024
This section presents the detailed Pearson Correlation Coefficient (PCC) results for cross-subject evaluations, complementing the MSE metrics discussed in the main text. Examining Tables 9 and 10, we observe a significant decrease in cross-subject PCC values compared to the within-subject PCC results presented in Tables 1 and 3. This performance degradation aligns with our findings in Section 4.4 and Section A.3, which demonstrate that during object recognition tasks, brain signals from different subjects exhibit substantial variations in spatial extent and amplitude magnitudes, even when response locations remain relatively consistent. These inter-subject variations pose significant challenges for developing models that can generalize effectively across individuals.
A.3 Additional Topography Visualization

### v2

@@ Abstract
Visual prostheses have shown great potential in restoring vision for blind individuals. However, while researchers have successfully utilized M/EEG signals to evoke visual perceptions during the brain decoding stage of visual prostheses, the complementary process—converting images to M/EEG signals in the brain encoding stage—remains largely unexplored. Thus, we present the first image-to-brain signal (M/EEG) framework based on denoising diffusion probabilistic models enhanced with cross-attention mechanisms. Our framework consists of two key architectural components: a pre-trained CLIP visual encoder that extracts rich semantic representations from input images, and a cross-attention enhanced U-Net diffusion model that learns to reconstruct biologically plausible brain signals through iterative denoising. Unlike conventional generative models that rely on simple concatenation for conditioning, our cross-attention modules enable dynamic interaction between visual features and brain signal representations, facilitating fine-grained alignment during the generation process. Furthermore, we evaluate our framework on two multimodal datasets (THINGS-EEG2 and THINGS-MEG) to demonstrate its effectiveness in generating biologically plausible brain signals. Additionally, we pioneer the visualization of M/EEG topographies across all subjects in both datasets, providing intuitive demonstrations of intra-subject and inter-subject variations in brain signals.
@@ 1 Introduction
Visual prostheses are advanced medical devices designed to restore partial vision for individuals with severe visual impairments or blindness, often caused by conditions such as retinitis pigmentosa (RP) and age-related macular degeneration (AMD) (Zrenner, 2013; Busskamp & Roska, 2011). These devices use an image sensor to capture external visual scenes and a processing framework to predict stimuli for an implanted electrode array (Humayun et al., 2012; Goetz & Palanker, 2016; Soltan et al., 2018) (we call this process brain encoding). The electrode array stimulates ganglion cells with the predicted stimuli, evoking visual perception (a pattern of localized light flashes, ’visual percept’, or ’phosphene’) in the retina (van der Grinten et al., 2024; Blom et al., 2010; Berry et al., 2017; Sahel et al., 2021; Granley et al., 2023) (this process is also referred to as brain decoding (Benchetrit et al., 2023)). The framework of visual prostheses is illustrated in Figure 1.
In the past few years, brain decoding has made significant progress (Lin et al., 2022b; Scotti et al., 2023; Wang et al., 2024b; Li et al., 2024). Specifically, Mind-Reader (Lin et al., 2022b), MindEye (Scotti et al., 2023), and MindBridge (Wang et al., 2024b) utilize the high spatial resolution of functional magnetic resonance imaging (fMRI) to generate phosphenes. However, due to the high cost and low temporal resolution of fMRI limiting their applications in brain-computer interfaces (BCIs), Li et al. (Li et al., 2024) not only leverage the high temporal resolution of electroencephalography (EEG) signals to evoke visual percepts, but also demonstrate the versatility of their work on magnetoencephalography (MEG) signals. More importantly, these studies (Lin et al., 2022b; Scotti et al., 2023; Wang et al., 2024b; Li et al., 2024) utilize multimodal datasets (Allen et al., 2022; Gifford et al., 2022) that include not only brain signals but also image data. Therefore, when training models, whether brain signals or image data are required, corresponding supervised signals can be provided to validate the model’s output.
@@
Figure 1: The framework of the visual prostheses. Visual prostheses utilize an image sensor to capture natural scenes. A processing framework takes the recorded signals as input and predicts the stimuli for the retinal prosthesis. A phosphene model receives stimulation from the implanted prosthesis and evokes visual perception (or ’phosphene’). The performance of the framework is evaluated by comparing the similarity between the original image and the visual perception.
To address the aforementioned issues, we propose an innovative image-to-brain framework that for the first time achieves the conversion of images to M/EEG signals. We employ a denoising diffusion probabilistic model (DDPM) (Ho et al., 2020) enhanced with cross-attention mechanisms. This framework consists of two core components: a CLIP visual encoder and a cross-attention enhanced U-Net diffusion model. The CLIP visual encoder extracts rich semantic representations from input images using pre-trained Vision Transformer (ViT-L/14) (Radford et al., 2021). The U-Net diffusion model reconstructs brain signals through iterative denoising, while cross-attention mechanisms enable fine-grained alignment between visual features and brain signal representations during the generation process. To validate our method’s effectiveness, we conduct experiments on two multimodal datasets (THINGS-EEG2 and THINGS-MEG) containing both brain signals and image data. With these datasets, we can directly learn the mapping from images to brain signals using the ground truth brain responses as supervision signals.
@@
  • We propose the first image-to-brain signal (M/EEG) framework based on diffusion models that achieves conversion from images to brain signals, advancing the technical foundation for visual prostheses.
@@
  • We validate our method through comprehensive experiments and plot M/EEG topographies for each subject on both datasets to intuitively demonstrate the intra-subject variations and inter-subject variations of M/EEG signals.
@@ 2 Related Works
In recent years, brain decoding has made significant progress by leveraging the powerful generative capabilities of diffusion models (Lin et al., 2022b; Scotti et al., 2023; Wang et al., 2024b; Li et al., 2024; Xu et al., 2023). In contrast, the development of brain encoding has progressed relatively slowly. Despite ongoing research efforts to improve the quality of predicted stimuli (Granley et al., 2022; 2023; Wang et al., 2024a), these studies fail to utilize real stimuli as supervised signals for evaluating the biological similarity of predicted stimuli, thereby limiting the vision restoration efficacy of visual prostheses to a low level (Montazeri et al., 2019).
To address the issues mentioned above, we use brain signals (M/EEG) from multimodal datasets (THINGS-EEG2, THINGS-MEG) as supervised signals to improve the biological similarity of predicted stimuli, thereby refining the image-to-brain framework.
@@ 3.3.2 Cross-Attention Mechanism
The key innovation of our approach lies in the integration of cross-attention mechanisms within the U-Net architecture. Unlike conventional generative models that use simple concatenation or addition for conditioning, our cross-attention modules (Lin et al., 2022a) enable fine-grained alignment between visual features and brain signals. We modify the standard U-Net by incorporating cross-attention blocks in both the encoder and decoder paths. These blocks enable dynamic interaction between the brain signal representations and visual features during the denoising process. For each cross-attention layer, given the intermediate brain signal representation $\mathbf{H}_{\text{brain}}$ and visual embedding $\mathbf{h}_{\text{img}}$, the cross-attention is computed as:
@@ 3.5 Test Stage
During testing, we start with pure Gaussian noise $\mathbf{y}_{T}\sim\mathcal{N}(0,\mathbf{I})$ and iteratively denoise it using the U-Net decoder $\mathcal{D}_{\text{dec}}$. Given an input image, we first extract visual features:
@@
Then, for each time step $t=T,T-1,\ldots,1$, we perform the denoising step using only the U-Net decoder $\mathcal{D}_{\text{dec}}$ with the image embeddings as conditioning:
$\mathbf{y}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{y}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\bm{\epsilon}_{\theta}^{\text{dec}}(\mathbf{y}_{t},t,\mathbf{h}_{\text{img}})\right)+\sigma_{t}\mathbf{z},$ (14)
where $\mathbf{z}\sim\mathcal{N}(0,\mathbf{I})$, $\sigma_{t}$ is the noise variance at step $t$, and $\bm{\epsilon}_{\theta}^{\text{dec}}$ represents the noise prediction computed using the U-Net decoder $\mathcal{D}_{\text{dec}}$ conditioned on the image embeddings $\mathbf{h}_{\text{img}}$. The final output $\mathbf{y}_{0}$ represents the generated brain signal corresponding to the input image.
@@ 4.2 Experiment Details
During training, we use the complete framework shown in Figure 2 with paired image-brain signal data. For testing, we perform image-to-brain signal conversion by first extracting image embeddings via the CLIP visual encoder, then using these embeddings as conditioning information to guide the U-Net decoder $\mathcal{D}_{\text{dec}}$ through iterative denoising from Gaussian noise to generate brain signals.
Our diffusion model employs a UNet2DConditionModel from the Hugging Face diffusers library (von Platen et al., 2022) with dataset-specific configurations. For EEG signals, we use a sample size of $(63,250)$ with 63 channels and 250 temporal sampling points. For MEG signals, we configure the model with a sample size of $(271,200)$ with 271 channels and 200 temporal sampling points. Both models use 1 input and output channel, 4 downsampling and upsampling blocks with channel dimensions of $(128,256,512,512)$. The downsampling path consists of: two DownBlock2D layers followed by two CrossAttnDownBlock2D layers, while the upsampling path includes: two CrossAttnUpBlock2D layers followed by two UpBlock2D layers. Cross-attention mechanisms are specifically integrated in the deeper layers to enable fine-grained alignment between visual features and brain signal representations (Ablation study on cross-modal learning can be seen in Appendix A.1). The cross-attention dimension is set to 768 to match the CLIP embedding dimension. We use the ViT-L/14 variant of CLIP as our visual encoder to extract rich semantic representations from input images.
@@ 4.3 Performance Evaluation
Table 3: Cross-subject PCC results on THINGS-EEG2 dataset
Train Subject Test Subject Source Stats
1 2 3 4 5 6 7 8 9 10 Mean Std
1 0.145 0.132 0.152 0.128 0.106 0.161 0.151 0.084 0.163 0.136 0.025
2 0.107 0.078 0.141 0.088 0.096 0.099 0.125 0.072 0.122 0.103 0.024
3 0.115 0.096 0.120 0.125 0.086 0.070 0.096 0.053 0.134 0.099 0.027
4 0.112 0.121 0.095 0.110 0.087 0.113 0.125 0.094 0.108 0.107 0.012
5 0.079 0.087 0.082 0.105 0.080 0.071 0.111 0.058 0.090 0.084 0.017
6 0.066 0.077 0.059 0.098 0.091 0.078 0.106 0.078 0.098 0.083 0.016
7 0.123 0.099 0.051 0.134 0.091 0.086 0.106 0.108 0.094 0.099 0.024
8 0.110 0.129 0.089 0.142 0.137 0.131 0.119 0.076 0.125 0.117 0.023
9 0.067 0.081 0.038 0.107 0.071 0.074 0.101 0.069 0.070 0.075 0.020
10 0.140 0.146 0.140 0.132 0.122 0.119 0.112 0.132 0.077 0.124 0.021
@@
Table 6: Cross-subject PCC results on THINGS-MEG dataset
Train Subject Test Subject Source Stats
1 2 3 4 Mean Std
1 0.080 0.038 0.069 0.062 0.021
2 0.088 0.075 0.051 0.071 0.019
3 0.055 0.071 0.045 0.057 0.013
4 0.027 0.033 0.020 0.027 0.007
Target Mean 0.057 0.061 0.044 0.055 0.054
Target Std 0.031 0.024 0.028 0.012 0.024
@@
Within-subject Performance: Tables 1 and 4 present the within-subject results, where models are trained and tested on data from the same subject. For the THINGS-EEG2 dataset, our method achieves an average MSE of 0.217 and PCC of 0.188 across 10 subjects. The performance varies across subjects, with Subject 1 achieving the best MSE (0.178) and Subject 8 showing the highest PCC (0.231). For the THINGS-MEG dataset with 4 subjects, we obtain an average MSE of 0.763 and PCC of 0.122. Among all subjects, Subject 1 achieves the lowest MSE of 0.607, while Subject 2 obtains the highest PCC of 0.198.
Cross-subject Generalization: From Tables 2 and 5, we observe that the cross-subject MSE averages (0.231 for EEG, 0.880 for MEG) are higher than the within-subject averages (0.217 for EEG, 0.763 for MEG), confirming the performance degradation in cross-subject scenarios. Examining cross-subject PCC results in Tables 3 and 6, we observe a significant decrease in cross-subject PCC values compared to the within-subject PCC results presented in Tables 1 and 4. These performance degradations align with our findings in Section 4.4 and Appendix A.2, which demonstrate that during object recognition tasks, brain signals from different subjects exhibit substantial variations in spatial extent and amplitude magnitudes, even when response locations remain relatively consistent. These inter-subject (cross-subject) variations pose significant challenges for developing models that can generalize effectively across individuals.
@@ 4.4 Visualization Analysis
I have plotted the topography for all subjects in both the THINGS-EEG2 and THINGS-MEG datasets, as detailed in Figures 3, 4, 5, 6, 7, and 8 (See Appendix A.2 for Figures 5, 6, 7, and 8).
@@
5 Conclusion
In conclusion, we present the first image-to-brain framework (we call image-to-brain process as brain encoding). This framework uses diffusion models to complete the brain signal reconstruction task and uses cross-attention to achieve alignment between the two modalities of images and brain signals. We conduct experiments on both THINGS-EEG2 and THINGS-MEG datasets, demonstrating the compatibility of our framework with both EEG and MEG signals. Meanwhile, we also plot topographies of EEG and MEG signals, allowing us to more intuitively observe the conditions of these two datasets.
@@ A.1 Ablation Study on Cross-Modal Learning
Tables 7 and 9 present the MSE results for the THINGS-EEG2 and THINGS-MEG datasets, respectively. Our cross-attention approach achieves the lowest average MSE on both datasets (0.217 for EEG and 0.763 for MEG), demonstrating superior performance compared to concatenation (0.228 for EEG and 0.808 for MEG) and addition (0.227 for EEG and 0.811 for MEG) methods.
Tables 8 and 10 present the results on the THINGS-EEG2 and THINGS-MEG datasets, respectively, with all values reported as Pearson Correlation Coefficient (PCC). The cross-attention mechanism achieves the highest average PCC on both datasets (0.188 for EEG and 0.122 for MEG), outperforming both concatenation (0.180 for EEG and 0.109 for MEG) and addition (0.172 for EEG and 0.118 for MEG) methods. The consistent advantage of cross-attention across both datasets and across different evaluation metrics indicates that explicitly modeling the dynamic interactions between brain signals and visual features leads to better performance in cross-modal learning tasks.
A.2 Additional Topography Visualization